# Allocation Multiplicity: Evaluating the Promises of the Rashomon Set

### v1

@@ 5.1. Data and Methods
Least Discriminatory Allocation 0.25 0.32 0.86 1.12 1.12 1.19 1.18
0.25 0.55 0.92 1.10 1.12 1.18 1.20
0.50 0.32 1.21 1.26 1.25 1.31 1.28
0.50 0.55 0.90 1.25 1.24 1.30 1.29
Outcome Homogenization 0.25 0.32 0.80 0.95 0.97 0.99 0.99
0.25 0.55 0.71 0.92 0.95 0.98 0.99
0.50 0.32 0.69 0.95 0.96 0.99 0.98
0.50 0.55 0.76 0.94 0.96 0.99 0.99
Age Homogenization 0.25 0.32 2.49 2.35 2.33 2.36 2.36
0.25 0.55 2.56 2.20 2.21 2.31 2.33
0.50 0.32 2.52 2.51 2.48 2.47 2.47
0.50 0.55 2.55 2.49 2.46 2.45 2.45
Table 3. Evaluating the Promises of the Rashomon Set. The “least discriminatory” allocation is based on the threshold test metric, which evaluates the ratio of chronic illnesses between selected black and white patients. Outcome homogenization measures the pairwise consistency of outcomes for individuals. Age homogenization measures the entropy over the age distribution of selected patients, where a lower entropy indicates a more narrow and unrepresentative distribution.
5.2. Evaluating Promises of the Rashomon Set
For each Rashomon set sampling method, we compare its allocations to the full set of equal-utility allocations, revealing that the recovered allocations fall short of the Rashomon set’s potential.
In real-world healthcare allocations, Obermeyer et al. (2019) highlight the fact that selected black patients are considerably sicker than selected white patients in terms of how many active chronic illnesses they have. This indicates discrimination based on the threshold test, which evaluates if the ratio of active chronic illnesses between selected black and white patients is greater than 1. In addition to its unfairness, this allocation is concerning from a prioritarian point of view: there are many black patients who were not selected who are sicker than the white patients who were selected. Across recovered allocations from the Rashomon set, we consider the least discriminatory allocation to have the smallest ratio based on the threshold test. Figure 4(a) shows that the smallest ratio recovered by all 4 methods is around 1.2, indicating that this “least discriminatory allocation” is still discriminatory. In contrast, the space of all equal-utility allocations contains many allocations with threshold test ratios at or less than 1. We specifically compare to an ideal equal-utility allocation that first selects $k^{\prime}$ qualified patients in descending order by number of chronic illnesses, then selects $k-k^{\prime}$ unqualified patients in the same fashion, alternating between black and white patients when there are ties. As Figure 4(a) shows, this results in a least [^11] We consider this the “least discriminatory” equal-utility allocation because it is the least discriminatory according to the normative goal of selecting patients based on how sick they are. Strictly speaking, an equal-utility allocation with a lower ratio could be obtained by adversarially choosing the least-ill black patients and the most-ill white patients. discriminatory equal-utility allocation with a threshold test ratio around 0.9.
Even if decision-makers use different models from the Rashomon set, outcome homogenization remains high and some qualified individuals are still systemically rejected.
We compare the pairwise consistency of individual outcomes between recovered allocations from the Rashomon set and all equal-utility allocations. Pairwise consistency represents the probability an individual would receive the same outcome under two random allocations from the set under consideration (Cooper et al., 2024). As Table 3 shows, the pairwise consistency is around 95% for recovered allocations from the Rashomon set. This is much higher than the pairwise consistency across all equal-utility allocations, which is around 75%. If decision-makers exclusively choose models from the Rashomon set, the high pairwise consistency implies that many individuals may be systemically rejected, even if they are qualified. Figure 4(b) analyzes the selection probability for qualified individuals across recovered allocations from the Rashomon set. Depending on the method, between 15% to 30% of qualified individuals are never selected in any allocation from the Rashomon set.
Consider the scenario where a fairness-oriented decision-maker uses the Rashomon set to average predictions and then selects patients with the top $k$ highest predictions from this ensemble. We show that the selected individuals under this allocation have different features from the selected individuals under a randomly-drawn equal-utility allocation. We focus on a specific feature – the age of selected patients – because of its potential to be homogenized in healthcare allocations. A stark example is the UK’s liver transplant matching algorithm, which systematically denied transplants to younger patients regardless of medical urgency (Narayanan et al., 2024). Figure 4(c) shows a similar result for the ensemble allocation from the Rashomon set: the age of selected patients is skewed older. While older patients are expected to be sicker and more qualified, Figure 4(c) also shows the age distribution of selected patients under a randomly-drawn equal-utility allocation, which chooses significantly more younger patients. Table 3 measures this phenomenon using the Shannon entropy of age brackets across selected patients. The recovered allocations from the Rashomon set have a lower entropy, indicating more homogenization in the age of selected patients.
6. Discussion
(a) Many qualified individuals have “high-risk” or low prediction values, indicating structural biases or underspecification in their risk scores.
(b) The lowest-risk allocation and the 10,000-th lowest-risk allocation, along with the allocations in between, have utility within 0.01 of each other on average (shaded region is S.D.).
Figure 5. (a) Risk Distribution of Qualified Individuals; (b) Low Deviations in Utility for Many “Riskier” Allocations
Our results show that conventional methods of using the Rashomon set do not fulfill its promises in resource-constrained allocations. Since we have established that model multiplicity is only useful insofar as it yields allocation multiplicity, how can we find more equal-utility allocations? We discuss three approaches below as avenues for future research.
Improving Sampling from the Rashomon Set
Allocation multiplicity should serve as a new benchmark for empirical methods that aim to find Rashomon models, especially given that a growing area of work involves developing and improving such methods (Rudin et al., 2024; Black et al., 2022). As Table 2 shows, all the methods that we explore find many different models in the Rashomon set. However, some methods perform better than others in terms of how many unique allocations they recover under a top $k$ mapping. We find that using different subsets of features and training data recovers more unique allocations than shuffling data order or perturbing model weights. But given that the former methods involve re-training different models from scratch, there are computational limitations to their use in practice. Even when many unique allocations are recovered, we further show that they fall short of what the full-space of equal-utility allocations promises. We hope the perspective of allocation multiplicity inspires future work to develop methods of finding models that deliver on these promises.
Mitigating Structural Biases and Underspecification
Addressing structural biases and underspecification in how risk scores are generated may lead to equal-utility allocations that status-quo models fail to reflect. Even if we sample extensively from the Rashomon set, the features used to model risk may lack the richness needed to distinguish between high-risk and low-risk individuals, particularly within certain subgroups. For example, in healthcare allocation, models trained on datasets that underrepresent low-income populations may fail to account for systemic barriers, such as limited access to preventive care or higher exposure to environmental risks (Obermeyer et al., 2019). As a result, these individuals are often misclassified as high-risk, even when they might benefit greatly from treatment (c.f. Figure 5(a)). This deficiency is compounded by incomplete or biased training data that limits the ability of models to learn meaningful patterns for marginalized groups. Consequently, even diverse models from the Rashomon set can converge on overly simplistic or biased risk estimates, systematically excluding certain qualified individuals and narrowing the recovered allocation space.
Non-Deterministic Mappings Between Predictions and Decisions
Our case study focused on the limitations of allocations recovered by deterministically selecting individuals with the lowest risk. However, there are many stochastic procedures to map predictions to decisions (Jain et al., 2024a). Decision-makers could use a weighted lottery based on predicted risk scores, or randomize decisions for individuals near the top $k$ threshold. These methods may better explore the space of allocations than trying to find more models in the Rashomon set, which are not guaranteed to produce different allocations under the deterministic paradigm. While stochastic procedures may involve trade-offs with expected utility, there may be little or no trade-off with realized utility, especially given the exponential number of equal-utility allocations that exist (c.f. Figure 5(b)). We hope that the knowledge of allocation multiplicity inspires a paradigm shift in how decision-makers interpret and act on model predictions, emphasizing flexibility and equity over rigid, risk-based thresholds.

### v3

@@ Acknowledgements
Least Discriminatory Allocation $\dagger$ 0.10 0.19 0.90 $\pm$ 0.08 0.97 $\pm$ 0.14 1.00 $\pm$ 0.12 1.09 $\pm$ 0.12 1.10 $\pm$ 0.13
0.10 0.32 0.95 $\pm$ 0.09 0.92 $\pm$ 0.14 0.99 $\pm$ 0.13 1.08 $\pm$ 0.12 1.11 $\pm$ 0.13
0.10 0.55 1.00 $\pm$ 0.09 0.90 $\pm$ 0.14 0.99 $\pm$ 0.15 1.09 $\pm$ 0.13 1.13 $\pm$ 0.13
0.25 0.19 1.06 $\pm$ 0.09 1.12 $\pm$ 0.13 1.14 $\pm$ 0.12 1.19 $\pm$ 0.13 1.16 $\pm$ 0.12
0.25 0.32 0.86 $\pm$ 0.08 1.12 $\pm$ 0.13 1.12 $\pm$ 0.12 1.19 $\pm$ 0.13 1.18 $\pm$ 0.12
0.25 0.55 0.92 $\pm$ 0.08 1.10 $\pm$ 0.14 1.12 $\pm$ 0.13 1.18 $\pm$ 0.13 1.20 $\pm$ 0.13
0.50 0.19 1.20 $\pm$ 0.12 1.25 $\pm$ 0.14 1.25 $\pm$ 0.13 1.30 $\pm$ 0.14 1.28 $\pm$ 0.14
0.50 0.32 1.21 $\pm$ 0.12 1.26 $\pm$ 0.14 1.25 $\pm$ 0.13 1.31 $\pm$ 0.14 1.28 $\pm$ 0.14
0.50 0.55 0.90 $\pm$ 0.10 1.25 $\pm$ 0.14 1.24 $\pm$ 0.13 1.30 $\pm$ 0.14 1.29 $\pm$ 0.14
Outcome Homogenization $\ddagger$ 0.10 0.19 0.88 $\pm$ 0.01 0.96 $\pm$ 0.00 0.98 $\pm$ 0.00 0.99 $\pm$ 0.00 0.99 $\pm$ 0.00
0.10 0.32 0.85 $\pm$ 0.00 0.95 $\pm$ 0.00 0.97 $\pm$ 0.00 0.99 $\pm$ 0.00 1.00 $\pm$ 0.00
0.10 0.55 0.83 $\pm$ 0.00 0.94 $\pm$ 0.00 0.96 $\pm$ 0.00 0.99 $\pm$ 0.00 1.00 $\pm$ 0.00
0.25 0.19 0.81 $\pm$ 0.01 0.96 $\pm$ 0.01 0.97 $\pm$ 0.00 0.99 $\pm$ 0.00 0.98 $\pm$ 0.00
0.25 0.32 0.80 $\pm$ 0.01 0.95 $\pm$ 0.00 0.97 $\pm$ 0.00 0.99 $\pm$ 0.00 0.99 $\pm$ 0.00
0.25 0.55 0.71 $\pm$ 0.01 0.92 $\pm$ 0.01 0.95 $\pm$ 0.00 0.98 $\pm$ 0.00 0.99 $\pm$ 0.00
0.50 0.19 0.60 $\pm$ 0.01 0.94 $\pm$ 0.01 0.95 $\pm$ 0.01 0.99 $\pm$ 0.00 0.97 $\pm$ 0.00
0.50 0.32 0.69 $\pm$ 0.01 0.95 $\pm$ 0.01 0.96 $\pm$ 0.01 0.99 $\pm$ 0.00 0.98 $\pm$ 0.00
0.50 0.55 0.76 $\pm$ 0.02 0.94 $\pm$ 0.01 0.96 $\pm$ 0.01 0.99 $\pm$ 0.00 0.99 $\pm$ 0.00
Age Homogenization $\star$ 0.10 0.19 2.46 $\pm$ 0.05 2.19 $\pm$ 0.10 2.27 $\pm$ 0.08 2.30 $\pm$ 0.09 2.30 $\pm$ 0.09
0.10 0.32 2.49 $\pm$ 0.04 2.04 $\pm$ 0.10 2.09 $\pm$ 0.09 2.24 $\pm$ 0.10 2.28 $\pm$ 0.10
0.10 0.55 2.56 $\pm$ 0.03 1.82 $\pm$ 0.11 1.84 $\pm$ 0.11 2.24 $\pm$ 0.09 2.28 $\pm$ 0.10
0.25 0.19 2.50 $\pm$ 0.04 2.43 $\pm$ 0.05 2.42 $\pm$ 0.05 2.39 $\pm$ 0.05 2.40 $\pm$ 0.05
0.25 0.32 2.49 $\pm$ 0.04 2.35 $\pm$ 0.06 2.33 $\pm$ 0.05 2.36 $\pm$ 0.05 2.36 $\pm$ 0.05
0.25 0.55 2.56 $\pm$ 0.03 2.20 $\pm$ 0.05 2.21 $\pm$ 0.06 2.31 $\pm$ 0.05 2.33 $\pm$ 0.06
0.50 0.19 2.55 $\pm$ 0.03 2.53 $\pm$ 0.03 2.48 $\pm$ 0.03 2.46 $\pm$ 0.04 2.47 $\pm$ 0.04
0.50 0.32 2.52 $\pm$ 0.03 2.51 $\pm$ 0.03 2.48 $\pm$ 0.03 2.47 $\pm$ 0.04 2.47 $\pm$ 0.04
0.50 0.55 2.55 $\pm$ 0.03 2.49 $\pm$ 0.03 2.46 $\pm$ 0.04 2.45 $\pm$ 0.04 2.45 $\pm$ 0.04
$\dagger$ The least discriminatory allocation shows the lowest ratio of active chronic illnesses between selected Black & White patients.
$\ddagger$ Outcome homogenization shows the probability of an individual receiving the same outcome under 2 random allocations.
$\star$ Age homogenization shows the Shannon entropy of age brackets across selected patients ($\uparrow$ entropy = $\downarrow$ homogenization).
Table 7. Outcome Homogenization Across the Rashomon Set (c.f. Figure 4(b))
Proportion of Qualified Individuals $k/n$ $n^{\prime}/n$ Different Bootstrapped Shuffled Weight
(Mean $\pm$ SD) Feature Subsets Training Data Data Order Perturbation
Rejected by all Rashomon Models 0.10 0.19 0.182 $\pm$ 0.028 0.323 $\pm$ 0.034 0.509 $\pm$ 0.028 0.499 $\pm$ 0.035
0.10 0.32 0.313 $\pm$ 0.025 0.451 $\pm$ 0.028 0.645 $\pm$ 0.027 0.683 $\pm$ 0.024
0.10 0.55 0.462 $\pm$ 0.031 0.592 $\pm$ 0.022 0.775 $\pm$ 0.014 0.811 $\pm$ 0.007
0.25 0.19 0.050 $\pm$ 0.017 0.077 $\pm$ 0.018 0.119 $\pm$ 0.021 0.098 $\pm$ 0.020
0.25 0.32 0.103 $\pm$ 0.019 0.163 $\pm$ 0.021 0.285 $\pm$ 0.026 0.265 $\pm$ 0.029
0.25 0.55 0.217 $\pm$ 0.028 0.311 $\pm$ 0.020 0.506 $\pm$ 0.017 0.541 $\pm$ 0.015
0.50 0.19 0.010 $\pm$ 0.009 0.014 $\pm$ 0.009 0.025 $\pm$ 0.012 0.020 $\pm$ 0.011
0.50 0.32 0.018 $\pm$ 0.008 0.032 $\pm$ 0.010 0.046 $\pm$ 0.012 0.038 $\pm$ 0.011
0.50 0.55 0.044 $\pm$ 0.015 0.081 $\pm$ 0.011 0.154 $\pm$ 0.019 0.142 $\pm$ 0.023
Multiple Outcomes From Rashomon Models 0.10 0.19 0.797 $\pm$ 0.038 0.521 $\pm$ 0.040 0.085 $\pm$ 0.031 0.081 $\pm$ 0.032
0.10 0.32 0.678 $\pm$ 0.026 0.489 $\pm$ 0.031 0.117 $\pm$ 0.035 0.041 $\pm$ 0.023
0.10 0.55 0.535 $\pm$ 0.032 0.376 $\pm$ 0.024 0.085 $\pm$ 0.023 0.018 $\pm$ 0.008
0.25 0.19 0.683 $\pm$ 0.194 0.133 $\pm$ 0.025 0.015 $\pm$ 0.009 0.053 $\pm$ 0.018
0.25 0.32 0.797 $\pm$ 0.064 0.408 $\pm$ 0.041 0.054 $\pm$ 0.018 0.083 $\pm$ 0.024
0.25 0.55 0.731 $\pm$ 0.056 0.497 $\pm$ 0.024 0.115 $\pm$ 0.020 0.040 $\pm$ 0.014
0.50 0.19 0.477 $\pm$ 0.144 0.057 $\pm$ 0.017 0.005 $\pm$ 0.005 0.015 $\pm$ 0.009
0.50 0.32 0.488 $\pm$ 0.140 0.075 $\pm$ 0.015 0.006 $\pm$ 0.005 0.023 $\pm$ 0.008
0.50 0.55 0.660 $\pm$ 0.079 0.307 $\pm$ 0.035 0.032 $\pm$ 0.011 0.048 $\pm$ 0.020
Accepted by all Rashomon Models 0.10 0.19 0.021 $\pm$ 0.022 0.157 $\pm$ 0.038 0.406 $\pm$ 0.038 0.419 $\pm$ 0.031
0.10 0.32 0.010 $\pm$ 0.007 0.061 $\pm$ 0.018 0.239 $\pm$ 0.023 0.277 $\pm$ 0.015
0.10 0.55 0.003 $\pm$ 0.003 0.031 $\pm$ 0.008 0.140 $\pm$ 0.012 0.171 $\pm$ 0.006
0.25 0.19 0.267 $\pm$ 0.189 0.790 $\pm$ 0.028 0.866 $\pm$ 0.022 0.849 $\pm$ 0.024
0.25 0.32 0.101 $\pm$ 0.053 0.429 $\pm$ 0.044 0.661 $\pm$ 0.033 0.651 $\pm$ 0.028
0.25 0.55 0.052 $\pm$ 0.033 0.192 $\pm$ 0.016 0.379 $\pm$ 0.015 0.419 $\pm$ 0.013
0.50 0.19 0.512 $\pm$ 0.140 0.929 $\pm$ 0.018 0.971 $\pm$ 0.013 0.966 $\pm$ 0.014
0.50 0.32 0.494 $\pm$ 0.136 0.894 $\pm$ 0.017 0.948 $\pm$ 0.013 0.939 $\pm$ 0.013
0.50 0.55 0.296 $\pm$ 0.067 0.613 $\pm$ 0.033 0.815 $\pm$ 0.019 0.810 $\pm$ 0.017
Table 8. Evaluating Stochastic Mappings Between Predictions & Decisions
Metric (Mean $\pm$ SD) $k/n$ $n^{\prime}/n$ All Equal-Utility Avg. Recovered Decision-Boundary Sigmoid-Logit
Allocations Rashomon Set (0.25kk, 0.50kk) (0.50kk, 1.0kk) (v=v=2) (v=v=5)
Allocation Utility $($$k^{\prime}/k$) 0.10 0.19 - 0.849 $\pm$ 0.033 0.850 $\pm$ 0.031 0.839 $\pm$ 0.030 0.832 $\pm$ 0.028 0.851 $\pm$ 0.032
0.10 0.32 - 0.929 $\pm$ 0.024 0.929 $\pm$ 0.026 0.924 $\pm$ 0.023 0.893 $\pm$ 0.019 0.917 $\pm$ 0.022
0.10 0.55 - 0.974 $\pm$ 0.014 0.976 $\pm$ 0.013 0.974 $\pm$ 0.012 0.926 $\pm$ 0.011 0.940 $\pm$ 0.012
0.25 0.19 - 0.658 $\pm$ 0.039 0.649 $\pm$ 0.034 0.623 $\pm$ 0.031 0.653 $\pm$ 0.038 0.658 $\pm$ 0.039
0.25 0.32 - 0.876 $\pm$ 0.021 0.869 $\pm$ 0.020 0.834 $\pm$ 0.021 0.851 $\pm$ 0.019 0.866 $\pm$ 0.019
0.25 0.55 - 0.952 $\pm$ 0.012 0.952 $\pm$ 0.012 0.943 $\pm$ 0.011 0.898 $\pm$ 0.012 0.898 $\pm$ 0.013
0.50 0.19 - 0.367 $\pm$ 0.023 0.365 $\pm$ 0.023 0.363 $\pm$ 0.023 0.364 $\pm$ 0.023 0.367 $\pm$ 0.023
0.50 0.32 - 0.604 $\pm$ 0.026 0.599 $\pm$ 0.025 0.593 $\pm$ 0.025 0.601 $\pm$ 0.025 0.603 $\pm$ 0.026
0.50 0.55 - 0.904 $\pm$ 0.013 0.885 $\pm$ 0.013 0.850 $\pm$ 0.013 0.868 $\pm$ 0.013 0.888 $\pm$ 0.013
Least Discriminatory Allocation $\dagger$ 0.10 0.19 0.90 $\pm$ 0.08 1.04 $\pm$ 0.13 1.04 $\pm$ 0.12 1.00 $\pm$ 0.13 0.94 $\pm$ 0.12 1.02 $\pm$ 0.12
0.10 0.32 0.95 $\pm$ 0.09 1.03 $\pm$ 0.13 1.07 $\pm$ 0.13 1.02 $\pm$ 0.14 0.78 $\pm$ 0.12 0.87 $\pm$ 0.12
0.10 0.55 1.00 $\pm$ 0.09 1.03 $\pm$ 0.14 1.09 $\pm$ 0.16 1.02 $\pm$ 0.16 0.61 $\pm$ 0.12 0.66 $\pm$ 0.11
0.25 0.19 1.06 $\pm$ 0.09 1.15 $\pm$ 0.12 1.10 $\pm$ 0.11 1.08 $\pm$ 0.11 1.09 $\pm$ 0.11 1.13 $\pm$ 0.12
0.25 0.32 0.86 $\pm$ 0.08 1.15 $\pm$ 0.12 1.11 $\pm$ 0.11 1.09 $\pm$ 0.12 1.02 $\pm$ 0.11 1.04 $\pm$ 0.11
0.25 0.55 0.92 $\pm$ 0.08 1.15 $\pm$ 0.13 1.14 $\pm$ 0.13 1.11 $\pm$ 0.13 0.87 $\pm$ 0.11 0.87 $\pm$ 0.12
0.50 0.19 1.20 $\pm$ 0.12 1.27 $\pm$ 0.14 1.24 $\pm$ 0.13 1.23 $\pm$ 0.13 1.22 $\pm$ 0.13 1.25 $\pm$ 0.14
0.50 0.32 1.21 $\pm$ 0.12 1.27 $\pm$ 0.14 1.23 $\pm$ 0.13 1.22 $\pm$ 0.13 1.23 $\pm$ 0.13 1.25 $\pm$ 0.13
0.50 0.55 0.90 $\pm$ 0.10 1.27 $\pm$ 0.14 1.23 $\pm$ 0.13 1.23 $\pm$ 0.14 1.13 $\pm$ 0.13 1.14 $\pm$ 0.13
Outcome Homogenization $\ddagger$ 0.10 0.19 0.88 $\pm$ 0.01 0.98 $\pm$ 0.00 0.97 $\pm$ 0.00 0.95 $\pm$ 0.00 0.94 $\pm$ 0.01 0.97 $\pm$ 0.00
0.10 0.32 0.85 $\pm$ 0.00 0.98 $\pm$ 0.00 0.97 $\pm$ 0.00 0.95 $\pm$ 0.00 0.89 $\pm$ 0.00 0.93 $\pm$ 0.01
0.10 0.55 0.83 $\pm$ 0.00 0.97 $\pm$ 0.00 0.97 $\pm$ 0.00 0.95 $\pm$ 0.00 0.85 $\pm$ 0.00 0.86 $\pm$ 0.00
0.25 0.19 0.81 $\pm$ 0.01 0.98 $\pm$ 0.00 0.94 $\pm$ 0.00 0.90 $\pm$ 0.00 0.94 $\pm$ 0.01 0.97 $\pm$ 0.00
0.25 0.32 0.80 $\pm$ 0.01 0.98 $\pm$ 0.00 0.94 $\pm$ 0.00 0.88 $\pm$ 0.00 0.89 $\pm$ 0.01 0.93 $\pm$ 0.02
0.25 0.55 0.71 $\pm$ 0.01 0.96 $\pm$ 0.00 0.94 $\pm$ 0.00 0.88 $\pm$ 0.00 0.73 $\pm$ 0.01 0.73 $\pm$ 0.01
0.50 0.19 0.60 $\pm$ 0.01 0.96 $\pm$ 0.01 0.89 $\pm$ 0.00 0.84 $\pm$ 0.01 0.85 $\pm$ 0.01 0.95 $\pm$ 0.01
0.50 0.32 0.69 $\pm$ 0.01 0.97 $\pm$ 0.01 0.89 $\pm$ 0.00 0.84 $\pm$ 0.01 0.90 $\pm$ 0.01 0.95 $\pm$ 0.01
0.50 0.55 0.76 $\pm$ 0.02 0.97 $\pm$ 0.01 0.89 $\pm$ 0.00 0.81 $\pm$ 0.01 0.84 $\pm$ 0.02 0.89 $\pm$ 0.02
Age Homogenization $\star$ 0.10 0.19 2.46 $\pm$ 0.05 2.26 $\pm$ 0.09 2.28 $\pm$ 0.09 2.30 $\pm$ 0.08 2.32 $\pm$ 0.07 2.27 $\pm$ 0.09
0.10 0.32 2.49 $\pm$ 0.04 2.16 $\pm$ 0.10 2.10 $\pm$ 0.11 2.14 $\pm$ 0.09 2.27 $\pm$ 0.06 2.17 $\pm$ 0.08
0.10 0.55 2.56 $\pm$ 0.03 2.04 $\pm$ 0.10 1.83 $\pm$ 0.13 1.92 $\pm$ 0.10 2.36 $\pm$ 0.03 2.27 $\pm$ 0.05
0.25 0.19 2.50 $\pm$ 0.04 2.41 $\pm$ 0.05 2.43 $\pm$ 0.05 2.44 $\pm$ 0.04 2.44 $\pm$ 0.05 2.43 $\pm$ 0.05
0.25 0.32 2.49 $\pm$ 0.04 2.35 $\pm$ 0.05 2.38 $\pm$ 0.05 2.43 $\pm$ 0.04 2.42 $\pm$ 0.04 2.37 $\pm$ 0.05
0.25 0.55 2.56 $\pm$ 0.03 2.26 $\pm$ 0.06 2.23 $\pm$ 0.06 2.27 $\pm$ 0.05 2.50 $\pm$ 0.03 2.49 $\pm$ 0.03
0.50 0.19 2.55 $\pm$ 0.03 2.48 $\pm$ 0.04 2.51 $\pm$ 0.03 2.51 $\pm$ 0.04 2.53 $\pm$ 0.03 2.48 $\pm$ 0.04
0.50 0.32 2.52 $\pm$ 0.03 2.48 $\pm$ 0.04 2.52 $\pm$ 0.03 2.52 $\pm$ 0.03 2.52 $\pm$ 0.03 2.50 $\pm$ 0.04
0.50 0.55 2.55 $\pm$ 0.03 2.46 $\pm$ 0.04 2.50 $\pm$ 0.03 2.52 $\pm$ 0.03 2.56 $\pm$ 0.03 2.56 $\pm$ 0.03
$\dagger$ The least discriminatory allocation shows the lowest ratio of active chronic illnesses between selected Black & White patients.
$\ddagger$ Outcome homogenization shows the probability of an individual receiving the same outcome under 2 random allocations.
$\star$ Age homogenization shows the Shannon entropy of age brackets across selected patients ($\uparrow$ entropy = $\downarrow$ homogenization).